2x+5y+3z=2


x+2y+2z=4


x+y+4z=11. Use the elementary row reduction to solve for the linear system.

1
Expert's answer
2022-04-05T14:43:38-0400

"\begin{pmatrix}

 2 & 5&3&|&2\\

 1 & 2&2&|&4\\1&1&4&|&11

\end{pmatrix}"



"R_{1}\leftrightarrow R_{3}\begin{pmatrix}

 1&1&4&|&11\\

 1 & 2&2&|&4\\2 & 5&3&|&2

\end{pmatrix}"



"-R_{1}+R_{2}\rightarrow R_{2}"

"-2R_{1}+R_{3}\rightarrow R_{3}" "\begin{pmatrix}
 1&1&4&|&11\\
 0 & 1&-2&|&-7\\0 & 3&-5&|&-20
\end{pmatrix}"



"-3R_{2}+R_{3}\rightarrow R_{3}" "\begin{pmatrix}
 1&1&4&|&11\\
 0 & 1&-2&|&-7\\0 & 0&1&|&1
\end{pmatrix}"




Hence, we have;


"x+y+4z=11"

"y-2z=-7"

"z=1"


"=> y=-7+2z=-7+2(1)=-5"

"=>x=11-4z-y=11-4(1)-(-5)=12"



"\u2234 x=12, y=-5, z=1"